OpenStudy (anonymous):

help please find f(f^-1(-8)), where f(x)=x^3 - 6

OpenStudy (shamil98):

Find the inverse of f(x), first.

OpenStudy (shamil98):

y = x^3 - 6 x = y^3 - 6 solve for y.

OpenStudy (helder_edwin):

actually, from the definition of the inverse of a function \[\large f(f^{-1}(-8))=(f\circ f^{-1})(-8)=-8 \]

OpenStudy (tkhunny):

Why find the inverse? The inverse is also a 1-to-1 function with no confusion for Domain or Range. \(f(f^{-1}(-8)) = -8\)

OpenStudy (anonymous):

so f(f^-1(-8))= what

OpenStudy (tkhunny):

We DO have to be careful first to notice that the inverse is ALSO a function WITHOUT any restriction on Domain or Range.

OpenStudy (shamil98):

oh mb.

OpenStudy (anonymous):

okay got it thanks